OpenStudy (anonymous):

please help! MEDAL REWARDED!!! What is the value of y in the solution to the following system of equations? 5x − 3y = −3 2x − 6y = −6

OpenStudy (texaschic101):

2x - 6y = -6 reduces to x - 3y = -3 now we have : 5x - 3y = -3 x - 3y = -3 --->(-5) ------------------ 5x - 3y = -3 -5x + 15y = 15 (result of multiply by -5) -------------------add 12y = 12 y = 1 This problem could have been solved without reducing the equation.....I just like messing with smaller numbers rather then bigger ones